OpenStudy (anonymous):

A magnetic field of magnitude B = 0.25 T is directed perpendicular to a circular loop of wire that has a radius of 0.47 m. What is the magnetic flux through the wire? ***How can I show the work for this? Thank you!!:)

OpenStudy (michele_laino):

here I am

OpenStudy (anonymous):

hi!!

OpenStudy (michele_laino):

the situation described by your exercise, is: 

OpenStudy (michele_laino):

by definition, the requested flux, is given by the susbsequent formula: \[\Large \Phi = BS = \pi {R^2}B\] where S is the area enclosed by our loop, namely it is the area of a circumference: \[\Large S = \pi {R^2}\]

OpenStudy (michele_laino):

so, what is: \[\Large \Phi = \pi {R^2}B = 3.14 \times {\left( {0.47} \right)^2} \times 0.25 = ...Weber\]

OpenStudy (anonymous):

sorry, my computer just froze!! but yes, we get 0.1734065 ?

OpenStudy (michele_laino):

that's right! We can round off that result like this: \[\Large \Phi = 0.173\;{\text{Weber}}\]

OpenStudy (anonymous):

ohh yay!! so that is all for this problem? that is the magnetic flux? 0.173 weber? :O

OpenStudy (michele_laino):

yes! Our task is completed!

OpenStudy (anonymous):

yay!! thank you so much!! :D

OpenStudy (michele_laino):

Thank you!! :D